Find the plane that passes through the point (2,5,3) and is perpendicular to the plane 2x-3y+4z+7=0
1
Expert's answer
2020-05-05T20:56:49-0400

First we define the normal to the plane. In general for the plane defined by the equation A*x+B*y+C*z+D normal vector is n=(A,B,C). So in our case we have n=(2,-3,4).




Let's define equation for the plane that passes through the point (2,5,3) in general:


A1*2+A2*5+A3*3 + A4 = 0.


For a plane passing through the origin and parallel to the plane above we have:


A1*2+A2*5+A3*3 = 0. (1)

Normal vector for this plane is n1=(A1,A2,A3).




For perpendicular planes, their normals are also perpendicular. Two nonzero vectors are perpendicular when their scalar product is zero. So we have 2*A1-3*A2+4*A3 = 0. (2)



Considering (1) and (2) we obtain the following system of linear equations:


2*A1+5*A2+3*A3=0


2*A1-3*A2+4*A3=0




It can be seen that our underdetermined system of linear equations must have at least one solution because of the rank of its coefficient matrix is equal to the rank of its augmented matrix (Kronecker–Capelli theorem). Since in an underdetermined system this rank is necessarily less than the number of unknowns, there are indeed an infinitude of solutions.




Let's subtract the second equation from the first equation:


8*A2-A3=0 or A3=8*A2 (3)



Let's substitute A3 from (3) into our system of linear equations:


2*A1+29*A2=0


2*A1+29*A2=0




Assuming A2 = 1 we get A1=-29/2=-14.5.


Substituting A2 =1 in (3) we have A3 = 8.




Thus, the desired equation of the plane has the form:


-14.5*x+y+8*z+A4=0, where A4 is constant.


Any plane parallel to the specified will be perpendicular to the given plane.

Planes are parallel if their normal vectors are collinear. The normal vector is determined by the coefficients A1, A2, A3. So for any A4 we get plane perpendicular to the given plane.


For definiteness let's put A4=0 (You can choose another constant number). So we have equation:


-14.5*x+y+8*z=0.